1. Students were asked to describe their politics as “Liberal”, “Moderate”, or “Conservative”. Here are the results:

Liberal Moderate Conservative Female 35 Male 50 36 21

a) What percent of the class is male?

b) What percent of the class considers themselves to be “Conservative”?

c) What percent of males in the class consider themselves to be “Conservative“?

d) What percent of all students in the class are males who consider themselves to be “Conservative”?

Please show all work.
Answer #1

a) What percent of the class is male?

Male = 50 + 44 + 21 = 115

Female = 35 + 36 + 6 = 77

Total = 115 + 77

= 192

%Male = (115 / 192 )*100

= 59.90 %

b) Conservatives = 6 + 21

= 27

% Conservatives = (27 /192)*100

= 14.06%

c) Males = 115

Conservative Males = 21

% Conservative Males = (21/115)*100

= 18.26%

d)

Conservative Males = 21

Total = 192

% Conservative Males = (21/192)*100

= 10.94%